OpenStudy (anonymous):

Assume the triangle has the given measurements. Solve for the remaining sides and angles A=(π/6,a=17.4,b=19.6

OpenStudy (anonymous):

I know that B=178.4, but I don't know how to find C and c

OpenStudy (anonymous):

@mtbender74

OpenStudy (anonymous):

Do you know the law of sines?

OpenStudy (anonymous):

yeah. Also I calculated 178.4 and I believe it is right, but it could be wrong lol

OpenStudy (anonymous):

so using the law of sines, you should be able to get each other piece...start by getting angle B which will then get you angle C. Once you have angle C, you can find side c... So my approach would be \[\frac{ \sin A }{ a } = \frac{ \sin B}{b}\] solve that for B then \[C = 2\pi - A - B \] then \[\frac{\sin A}{a} = \frac{\sin C}{c}\] solve that for c.

OpenStudy (anonymous):

i assume that your angles are intended to be radians...that's why i used 2pi

OpenStudy (anonymous):

is B=178.4

OpenStudy (anonymous):

haven't worked it out...lemme check

OpenStudy (anonymous):

no...were you using degrees or radians for your sin calculations?

OpenStudy (anonymous):

I got 178.4 degrees

OpenStudy (anonymous):

however it is A=pi/6 so it should be radians I think

OpenStudy (anonymous):

3.11 radians

OpenStudy (anonymous):

exactly... lemme set it up for you. \[\frac{\sin A}{a} = \frac{\sin B}{b}\] \[\frac{\sin{\frac{\pi}{6}}}{17.4} = \frac{\sin B}{19.6}\] \[\frac{.5}{17.4} = \frac{\sin B}{19.6}\] make sure your calculator is in radians mode and find B...

OpenStudy (anonymous):

oh...not exactly on 3.11 radian...

OpenStudy (anonymous):

3.11366739

OpenStudy (anonymous):

hmmm...i'm getting .6

OpenStudy (anonymous):

pi/6 is not exactly .5 though I would rather use .52

OpenStudy (anonymous):

sin pi/6 = .5

OpenStudy (anonymous):

ah ok so .6 seems correct for B

OpenStudy (anonymous):

so now, you can find C = 2pi - A - B

OpenStudy (anonymous):

what is .6 is radians or do I just leave the answer as 0.6?

OpenStudy (anonymous):

it is 0.6 radians...

OpenStudy (anonymous):

1 radian is about 60 degrees

OpenStudy (anonymous):

does B=0.6 or sinB=0.6 because sinpi/6 is different than pi/6

OpenStudy (anonymous):

B = 0.6 radians...

OpenStudy (anonymous):

Ok got it. so C=5.16

OpenStudy (anonymous):

that's what i got too... so now you have \[\frac{\sin A}{a} = \frac{\sin C}{c}\] \[\frac{\sin \frac{\pi}{6}}{17.4}=\frac{\sin 5.16}{c}\] solve for c...

OpenStudy (anonymous):

c=19.8? @mtbender74

OpenStudy (anonymous):

actually...our C is wrong...it should be pi - A - B...my mistake...

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

2.02 = C? are you sure?

OpenStudy (anonymous):

yeah..was trying to figure out why it wasn't working

OpenStudy (anonymous):

and that makes more sense with the answer i just got... sin 2.02 = .9

OpenStudy (anonymous):

so what is c?

OpenStudy (anonymous):

\[\frac{\sin(pi/6)}{17.4}=\frac{\sin 2.02}{c}\] \[\frac{.5}{17.4}=\frac{.9}{c}\] solve for c

OpenStudy (anonymous):

c=31.32. Ok, so overall the answers are B=0.6, C=2.02, and c=31.32. Correct?

OpenStudy (anonymous):

@mtbender74

OpenStudy (anonymous):

matches close enough to what i got...and the numbers make sense...

OpenStudy (anonymous):

lol so it is right? what did you get that was different

OpenStudy (anonymous):

c = 31.35...but i used the value from the calculator before rounding...

OpenStudy (anonymous):

would my answer be graded wrong lol

OpenStudy (anonymous):

as long as you show that you used .9 in the calculation *i* wouldn't if i were your teacher...and I was a teacher once upon a time... :)

OpenStudy (anonymous):

got it, thanks. maybe you can help sometime later!

OpenStudy (anonymous):

if i'm around...